Thursday, July 5, 2012

Status update to Board and SSC

Report on pODI Status - July 5, 2012

We are approaching the end of laboratory integration and testing of pODI, and beginning to prepare for shipping, installation, and commissioning.  We have encountered the normal number of problems that only become evident when you reach this point in the development of an instrument as complex as this.  So far, no show stoppers.  Here is a summary.

All 13 of the detectors in the pODI focal plane are operating properly, as are the focus sensors.  We have run the entire focal plane in "static imaging" mode, in which the central 3 X 3 array (approximately 24 arcmin square) integrates with amps turned off while the outer 4 are left active to acquire guide star data.  There is no detectable leakage of the glows from the outer, active detectors to the central region.  We have measured read noise, full well capacity, dark current, linearity and these all seem normal.  We are able to read small regions around a point representing a guide star at 20 Hz.  We have not yet run the detectors in orthogonal transfer (OT) mode on the focal plane (though we did in the test Dewar).  This is primarily due to a firmware upgrade that IfA put into the Stargrasp controllers that "broke" this mode. We'll get a fix from them in the next few days.  The IfA Stargrasp team visited and helped us with the optimization of the detectors.  We ended up with a small punch list of remaining work that they will do before we close out the contract. 

The Dewar has performed well in holding vacuum and maintaining operating temperature.  There is a little work to do on connectors, which will take place after we complete the detector testing.

The ISP, which holds the filter arms and ADC prisms, has been tested apart from the Dewar.  We had a significant problem with the filter arms, which are driven by two sets of worms and gears, a small, high-speed set and a large, low-speed set.  We had to modify the small ones by encasing them in a gearbox filled with grease, and add shims between filter arms to position the large ones more correctly.  We have now completed testing in our environmental chamber down to 0 degrees centigrade, and all functions seem to work acceptably.

Software is a major effort at this point.  On the instrument side, work is divided between supporting the detector testing and developing the capabilities that we will need for the commissioning effort.  The user interface is starting to take shape.  On the pipeline, portal, and archive side, the team is beginning acceptance testing of the first set of functional pieces.  Not all the pieces are there yet, so I won't call it an integration test, but we are trying to develop a model for formal testing of the complete system.  There will be an integrated system to test at the end of August with all the functionality needed to support commissioning.

Work on the telescope to prepare for ODI installation is also proceeding.  Delays in the dome repair work and aluminization have caused us to postpone our shipping date by a few days, but we have enough contingency in the installation schedule that this should not impact our overall schedule.  Our current plan is:

July 10 - completion of testing
July 12 - readiness and safety review
July 17 - begin packing
July 20 - first trip to mountain (front corrector and assorted small items)
July 24 - second trip to mountain (two large crates: Dewar and ISP)
July 27 - begin installation
Aug 16 - start engineering verification

Although the nominal date for first opening the dome is August 16, we will open earlier if things go well, and if the weather makes it possible.  Stay tuned!


No comments:

Post a Comment